### v1

# Finely Stratified Rerandomization Designs

Abstract
We study estimation and inference on causal parameters under finely stratified rerandomization designs, which use baseline covariates to match units into groups (e.g. matched pairs), then rerandomize within-group treatment assignments until a balance criterion is satisfied. We show that finely stratified rerandomization does partially linear regression adjustment “by design,” providing nonparametric control over the covariates used for stratification, and linear control over the rerandomization covariates. We also introduce novel rerandomization criteria, allowing for nonlinear imbalance metrics and proposing a minimax scheme that optimizes the balance criterion using pilot data or prior information provided by the researcher. While the asymptotic distribution of generalized method of moments (GMM) estimators under stratified rerandomization is generically non-Gaussian, we show how to restore asymptotic normality using optimal ex-post linear adjustment. This allows us to provide simple asymptotically exact inference methods for superpopulation parameters, as well as efficient conservative inference methods for finite population parameters.
1 Introduction
Stratified randomization is commonly used to increase statistical precision in experimental research. [^2] For example, Cytrynbaum, (2023) reports a survey of 50 experimental papers in the AER and AEJ from 2018-2023, where 57% used some form of stratified randomization. Recent theoretical work (e.g. Bai et al., (2021)) has shown that fine stratification, which randomizes within small groups of units tightly matched on baseline covariate information, makes unadjusted estimators like difference of means semiparametrically efficient. [^3] See Cytrynbaum, (2024), Armstrong, (2022), and Bai et al., (2024) for more detailed discussion. In finite samples, however, the performance of such designs can deteriorate rapidly with the dimension of the stratification variables due to a curse of dimensionality in matching. [^4] Under regularity conditions, the convergence rate of finite sample variance to asymptotic variance is $O(n^{-2/(d+1)})$ for dimension $d$ covariates, see Cytrynbaum, (2024). This motivates the search for alternative designs that insist upon nonparametric balance for a few important covariates, but only attempt to balance linear functions of the remaining variables. In this paper, we propose finely stratified rerandomization designs, which first tightly match the units into groups using a small set of important covariates, then rerandomize within groups until a balance criterion on the remaining covariates is satisfied.
Our first contribution is to derive the asymptotic distribution of generalized method of moments (GMM) estimators under stratified rerandomization, allowing estimation and inference on generic causal parameters defined by moment equalities. We consider both superpopulation and finite population parameters, the latter of which may be more appropriate for experiments run in a convenience sample (Abadie et al., (2014)). As in previous work on rerandomization (e.g. Li et al., (2018)), the asymptotic distribution of GMM estimators is an independent sum of a normal and a truncated normal term. Modulo this residual truncated term, we show that the asymptotic variance of unadjusted estimation under stratified rerandomization is the same as that of semiparametrically adjusted GMM (e.g. Graham, (2011)) under an iid design. Intuitively, stratified rerandomization implements partially linear regression adjustment “by design.”
Our second contribution is to introduce several novel forms of rerandomization based on nonlinear balance criteria. For example, we allow acceptance or rejection based on the difference of covariate density estimates within each treatment arm, attempting to balance nonlinear features of the covariate distribution. Similarly, we propose a design that rerandomizes until a nonlinear estimate of the propensity score is approximately constant, effectively forcing the covariates to have no predictive power for treatment assignments. In both cases, these nonlinear rerandomization schemes are asymptotically equivalent to standard rerandomization based on a difference of covariate means, but with an implicit choice of covariates and acceptance region, which we characterize.
Our third contribution is to study optimization of the balance criterion itself. We suggest a novel minimax approach that allows the researcher to specify prior information about the relationship between covariates and outcomes, then rerandomizes until the worst case correlation consistent with this prior information is small. If the prior information set contains the truth, this design provides strong control over the variance of the truncated normal term in the asymptotic distribution. Building on this, we show that if the prior information set is a confidence region estimated from pilot data, then this minimax design bounds the truncated normal variance with high probability.
Our fourth contribution is to provide simple t-statistic based inference methods for general causal parameters under stratified rerandomization designs. To do this, we first characterize and provide a feasible implementation of the optimal ex-post linear adjustment for GMM estimation under stratified rerandomization. [^5] This extends recent work on optimal adjustment under pure stratified randomization for ATE estimation, e.g. see Cytrynbaum, (2023), Bai et al., (2023), or Liu and Yang, (2020). Crucially, optimal ex-post adjustment makes the asymptotic distribution insensitive to the rerandomization acceptance criterion, removing the truncated normal term from the limiting distribution and restoring asymptotic normality. For superpopulation parameters, our inference methods are asymptotically exact. For finite population parameters, our methods are asymptotically conservative, but still exploit the efficiency gains from both stratified rerandomization and ex-post optimal adjustment.
1.1 Related Literature
This paper builds on the literature on fine stratification in econometrics as well as the literature on rerandomization in statistics. Stratified randomization has a long history in statistics, see Cochran, (1977) for a survey. Recent work on fine stratification in econometrics includes Bai et al., (2021), Bai, (2022), Cytrynbaum, (2024), Armstrong, (2022), and Bai et al., (2024). Some important theoretical contributions to the literature on rerandomization include Morgan and Rubin, (2012) and Li et al., (2018), Wang et al., (2021), and Wang and Li, (2022). We build on both of these literatures, studying the consequence of rerandomizing treatments within data-adaptive fine strata. We show that finely stratified rerandomization does semiparametric (partially linear) regression adjustment “by design,” providing nonparametric control over a few important variables and linear control over the rest.
For our main asymptotic theory (Section 3), the most closely related previous work is Wang et al., (2021) and Bai et al., (2024). Wang et al., (2021) study estimation of the sample average treatment effect (SATE) under stratified rerandomization, with quadratic imbalance metrics based on the Mahalanobis norm. We study rerandomization within data-adaptive fine strata, providing asymptotic theory for generic superpopulation and finite population causal parameters defined by moment equalities. We also allow for essentially arbitrary rerandomization acceptance criteria, not necessarily based on quadratic forms. Bai et al., (2024) study estimation of superpopulation parameters defined by moment equalities under pure stratified randomization. We extend these results to stratified rerandomization as well as generic finite population parameters, providing “SATE-like” versions of the parameters in Bai et al., (2024). [^6] These parameters can be seen as causal versions of the conditional estimand defined in Abadie et al., (2014). In concurrent work, Wang and Li, (2024) study GMM estimation of univariate superpopulation parameters under stratified rerandomization with fixed, discrete strata. We study significantly more general forms of stratification and rerandomization criteria than considered in their work, allowing for both finite and superpopulation parameters of arbitrary fixed dimension.
For nonlinear rerandomization (Section 4), the closest related results are Ding and Zhao, (2024) and Li et al., (2021). Ding and Zhao, (2024) rerandomize based on the p-value of a logistic regression coefficient, while we rerandomize until a general smooth propensity estimate is close to constant. To the best of our knowledge, we present the first asymptotic theory for rerandomization based on the difference of nonlinear (e.g. density) estimates. For acceptance region optimization (Section 5), the closest related results are Schindl and Branson, (2024), who study the optimal choice of norm for quadratic rerandomization, while Liu et al., (2023) chooses a specific quadratic rerandomization using a Bayesian criterion, in both cases for rerandomization without stratification. We provide a novel minimax approach that accepts or rejects based on the value of a convex penalty function, tailored to prior information provided by the researcher. Our work on optimal adjustment (Section 6) extends recent work on adjustment for stratified designs, e.g. Liu and Yang, (2020), Cytrynbaum, (2023), Bai et al., (2023), to stratified rerandomization and GMM parameters. Finally our inference methods (Section 7) build on previous work by Abadie and Imbens, (2008), Bai et al., (2021), and Cytrynbaum, (2024). To the best of our knowledge we provide the first asymptotically exact inference for causal GMM parameters under stratified rerandomization, as well as conservative inference for their finite population analogues.
2 Framework and Designs
Consider data $W_{i}=(R_{i},S_{i}(1),S_{i}(0))$ with $(W_{i})_{i=1}^{n}\operatorname{\overset{iid}{\sim}}F$. The $S_{i}(d)\in\mathbb{R}^{d_{S}}$ denote potential outcome vectors for a binary treatment $d\in\{0,1\}$, while $R_{i}$ denote other pre-treatment variables, such as covariates. For treatment assignments $D_{i}\in\{0,1\}$, the realized outcome $S_{i}=S_{i}(D_{i})=D_{i}S_{i}(1)+(1-D_{i})S_{i}(0)$. In what follows, for any array $(a_{i})_{i=1}^{n}$ we denote $E_{n}[a_{i}]=n^{-1}\sum_{i=1}^{n}a_{i}$, with $\bar{a}_{1}=E_{n}[a_{i}D_{i}]/E_{n}[D_{i}]$ and $\bar{a}_{0}=E_{n}[a_{i}(1-D_{i})]/E_{n}[(1-D_{i})]$ Next, we define stratified rerandomization designs.
Definition 2.1 (Stratified Rerandomization).
Let treatment proportions $p=l/k$ and suppose that $n$ is divisible by $k$ for notational simplicity.
  1. (1)
    (Stratification). Partition the experimental units into $n/k$ disjoint groups $g$ with $\{1,\dots,n\}=\bigcup_{g}g$ disjointly and $|g|=k$. Let $\psi=\psi(R)$ with $\psi\in\mathbb{R}^{d_{\psi}}$ denote a vector of stratification variables, which may be continuous or discrete. Suppose the groups satisfy the homogeneity condition [^7] The matching condition in Equation 2.1 was introduced by Bai et al., (2021) for matched pairs randomization ($k=2$). See Bai, (2022) and Cytrynbaum, (2024) for generalizations.
    $\frac{1}{n}\sum_{g}\sum_{i,j\in g}|\psi_{i}-\psi_{j}|_{2}^{2}=o_{p}(1).$ (2.1)
    Require that the groups only depend on the stratification variables $\psi_{1:n}$ and data-independent randomness $\pi_{n}$, so that $g=g(\psi_{1:n},\pi_{n})$ for each $g$.
  2. (2)
    (Randomization). Independently for each $|g|=k$, draw treatment variables $(D_{i})_{i\in g}$ by setting $D_{i}=1$ for exactly $l$ out of $k$ units, uniformly at random.
  3. (3)
    (Check Balance). For rerandomization covariates $h=h(R)$, consider an imbalance metric $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})+o_{p}(1)$. [^8] In particular, we require $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})+o_{p}(1)$ under “pure” stratified randomization, the design in steps (1) and (2) only, studied e.g. in Cytrynbaum, (2024). We give several examples below. For an acceptance region $A\subseteq\mathbb{R}^{d_{h}}$, check if the balance criterion $\mathcal{I}_{n}\in A$ is satisfied. If so, accept $D_{1:n}$. If not, repeat from the beginning of (2).
Intuitively, steps (1) and (2) describe a data-adaptive “matched k-tuples” design, while step (2) rerandomizes within k-tuples until the balance criterion is satisfied. Equation 2.1 is a tight-matching condition, requiring that the groups are clustered locally in $\psi$ space. Cytrynbaum, (2024) provides algorithms to match units into groups that satisfy this condition for any fixed $k$.
Example 2.2 (Matched Pairs Rerandomization).
[…]
Example 2.3 (Stratification).
Stratification without rerandomization can be obtained by setting $A=\mathbb{R}^{d_{h}}$ in Definition 2.1. Treatment effect estimation under such designs was studied in Bai, (2022), Cytrynbaum, (2024), and Bai et al., (2024). Definition 2.1 allows for fine stratification (also known as matched k-tuples), with the number of data-dependent groups $g=g(\psi_{1:n},\pi_{n})$ growing with $n$. It also allows for coarse stratification with fixed strata $s\in\{1,\dots,m\}$ and fixed $m$, as in Bugni et al., (2018), which can be obtained in this framework by setting $\psi=s$ and matching units into groups at random within the strata $\{i:s_{i}=k\}$.
Example 2.4 (Complete Randomization).
For $p=l/k$, we say that $D_{1:n}$ are completely randomized with probability $p$ if $P(D_{1:n}=d_{1:n})=1/\binom{n}{np}$ for all $d_{1:n}$ with $\sum_{i}d_{i}=np$ [^10] For notational simplicity, we may assume that $n=lk$ for some $l\in\mathbb{N}$. If so, we denote $D_{1:n}\sim\operatorname{CR}(p)$. Cytrynbaum, (2024) shows that $\operatorname{CR}(p)$ randomization can be obtained by setting $\psi=1$ and $A=\mathbb{R}^{d_{h}}$ in Definition 2.1, matching units into groups at random.
Causal Estimands. Next, we introduce a generic family of causal estimands defined by moment equalities. Let $g(D,R,S,\theta)\in\mathbb{R}^{d_{g}}$ be a score function for generalized method of moments (GMM) estimation. Recall $W=(R,S(1),S(0))$ and define the projection $\phi(W,\theta)=E[g(D,R,S,\theta)|W]$, so that $E[\phi(W,\theta)]=0$ $\iff$ $E[g(D,R,S,\theta)]=0$.
Definition 2.5 (Causal Estimands).
The superpopulation estimand $\theta_{0}$ is the unique solution to $E[\phi(W,\theta)]=0$. The finite population estimand $\theta_{n}$ is the unique solution to $E_{n}[\phi(W_{i},\theta)]=0$.
In what follows, we study GMM estimation of both $\theta_{0}$ and $\theta_{n}$ under stratified rerandomization designs, showing an asymptotic equivalence between stratified rerandomization and partially linear covariate adjustment. In particular, this framework allows us to introduce several useful finite population estimands $\theta_{n}$ that do not appear to have been considered previously in the literature. Note that GMM estimation of $\theta_{0}$ under pure stratification was studied in Bai et al., (2024) for the exactly identified case. Our finite population parameter $\theta_{n}$ can be viewed as a causal version of the finite population estimand defined in Abadie et al., (2014). [^11] See also the related finite population estimands studied under iid sampling and assignment in Xu, (2021) and Kakehi and Otsu, (2024).
Example 2.6 (ATE).
Define the Horvitz-Thompson weights $H=\frac{D-p}{p-p^{2}}$ and let $g(D,Y,\theta)=HY-\theta$, so that $\phi(W,\theta)=E[HY|W]-\theta=Y(1)-Y(0)-\theta$. Then $\theta_{0}=E[Y(1)-Y(0)]=\operatorname{ATE}$, the average treatment effect, and $\theta_{n}=E_{n}[Y_{i}(1)-Y_{i}(0)]=\operatorname{SATE}$, the sample average treatment effect.
For a more interesting example, consider the best parametric predictor of treatment effect heterogeneity in experiments with noncompliance.
Example 2.7 (LATE Heterogeneity).
Let $D(z)$ be potential treatments for a binary instrument $z\in\{0,1\}$. Let $Y(d)$ be the potential outcomes, with realized outcome $Y=Y(D(Z))$. Suppose $D(1)\geq D(0)$, and define compliance indicator $C=\mathds{1}(D(1)>D(0))$, assuming $E[C]>0$. Imbens and Angrist, (1994) define the local average treatment effect $\operatorname{LATE}=E[Y(1)-Y(0)|C=1]$. Let $H=(Z-p)/(p-p^{2})$ and consider the score function $g(Z,D,Y,X,\theta)=(HY-HD\cdot f(X,\theta))\nabla_{\theta}f(X,\theta)$. Using standard LATE manipulations,
$\phi(W,\theta)=E[g(Z,D,Y,X,\theta)|W]=C\cdot(Y(1)-Y(0)-f(X,\theta))\nabla_{\theta}f(X,\theta).$
This is the first order condition of a treatment effect prediction problem in the complier population. In particular, for $\tau\equiv Y(1)-Y(0)$, the parameter $\theta_{0}$ is the best parametric predictor of treatment effects for compliers:
$\theta_{0}=\operatorname*{argmin}_{\theta}E[(\tau-f(X,\theta))^{2}|C=1].$
For example, if $Y$ is binary then $Y(1)-Y(0)\in\{-1,0,1\}$, so a scaled link function model $f(X,\theta)=2L(X^{\prime}\theta)-1$ may be appropriate. We can easily estimate marginal effects by adding $m(X_{i},\theta,\beta)=\beta-(\partial/\partial\theta^{\prime})f(X_{i},\theta)$ to the score function.
Example 2.8 (Finite Population Heterogeneity).
Continuing Example 2.7, note that for $\tau_{i}=Y_{i}(1)-Y_{i}(0)$ the corresponding finite population parameter is
$\theta_{n}=\operatorname*{argmin}_{\theta}E_{n}[(\tau_{i}-f(X_{i},\theta))^{2}|C_{i}=1].$ (2.2)
We can view $\theta_{n}$ as a “SATE-like” version of $\theta_{0}$, the best parametric predictor of treatment effects in the within-sample complier population. $\theta_{n}$ may be a more appropriate target for experiments run in a convenience sample. If $f(X,\theta)=X^{\prime}\theta$ linear, then $\theta_{n}=\operatorname*{argmin}_{\theta}E_{n}[(\tau_{i}-X_{i}^{\prime}\theta)^{2}|C_{i}=1]$ is the within-sample best linear predictor. In the case of perfect compliance $C_{i}=1$ for all $i$, this is $\theta_{n}=\operatorname*{argmin}_{\theta}E_{n}[(\tau_{i}-X_{i}^{\prime}\theta)^{2}]$, a finite-sample version of the best linear predictor of the conditional average treatment effect (CATE). The case $X=1$ recovers $\theta_{n}=E_{n}[Y_{i}(1)-Y_{i}(0)|C_{i}=1]$, the finite-population LATE, studied e.g. in Ren, (2023). Our inference methods in Section 7 produce tighter confidence intervals for these finite population parameters than $\theta_{0}$, since we only need to account for the uncertainty due to random assignment, with no sampling uncertainty.
[…]
$\widehat{\theta}=\operatorname*{argmin}_{\theta\in\Theta}\widehat{g}(\theta)^{\prime}M_{n}^{\prime}\widehat{g}(\theta).$ (2.3)
In the exactly identified case, $\widehat{\theta}$ solves $\widehat{g}(\widehat{\theta})=0$. In the next section, we study generalized method of moments (GMM) estimation of the causal parameters $\theta_{0}$ and $\theta_{n}$ under stratified rerandomization.
3 Asymptotics for GMM Estimation
In this section, we characterize the asymptotic distribution of the GMM estimator $\widehat{\theta}$ under stratified rerandomization designs, as in Definition 2.1. We show that the variance under stratified rerandomization is proportional to the residuals of a partially linear regression model, up to a rerandomization imbalance term. In this sense, stratified rerandomization does partially linear regression adjustment “by design.” First, we state some technical regularity conditions that are needed for the following results.
Assumption 3.1 (Acceptance Region).
Suppose $A\in\mathbb{R}^{d_{h}}$ has non-empty interior and $\operatorname{Leb}(\partial A)=0$, [^13] Note that $\partial A$ denotes the boundary of $A$, the limit points of both $A$ and $A^{c}$. and require $E[\operatorname{Var}(h|\psi)]\succ 0$ and $E[|\psi|_{2}^{2}+|h|_{2}^{2}]<\infty$.
Next we state the technical conditions needed for GMM estimation. Define the matrix $G=E[(\partial/\partial\theta^{\prime})\phi(W,\theta)]|_{\theta=\theta_{0}}\in\mathbb{R}^{d_{g}\times d_{\theta}}$ and let $g_{d}(W,\theta)=g(d,R,S(d),\theta)$ for $d\in\{0,1\}$. Recall that $|B|_{F}^{2}=\sum_{ij}B_{ij}^{2}$ for any matrix $B$.
Assumption 3.2 (GMM).
The following conditions hold for $d\in\{0,1\}$:
  1. (a)
    We have $E[g_{d}(W,\theta_{0})^{2}]<\infty$ and $E[\sup_{\theta\in\Theta}|g_{d}(W,\theta)|_{2}]<\infty$. Also $\theta\to g_{d}(W,\theta)$ is continuous almost surely, and $\Theta$ is compact with $\theta_{0}$ in the interior.
  2. (b)
    The matrix $G$ is full rank, and $g_{0}(\theta)=0$ iff $\theta=\theta_{0}$.
  3. (c)
    There exists a neighborhood $\theta_{0}\in U\subseteq\Theta$ such that $G_{d}(W,\theta)\equiv\partial/\partial\theta^{\prime}g_{d}(W,\theta)$ exists and is continuous. Also $E[\sup_{\theta\in U}|\partial/\partial\theta^{\prime}g_{d}(W,\theta)|_{F}]<\infty$.
Compactness could likely be relaxed using concavity assumptions or a VC class condition, but we do not pursue this here. In what follows it will be conceptually useful to reparameterize the score function.
Orthogonal Expansion. Recall $\phi(W,\theta)=E[g(D,R,S,\theta)|W]$ for $W=(R,S(1),S(0))$. Define the assignment influence component $\delta(W,\theta)\equiv\operatorname{Var}(D)(g_{1}(W,\theta)-g_{0}(W,\theta))$. For the Horvitz-Thompson weights $H=(D-p)/(p-p^{2})$, a simple calculation shows that we can expand
$g(D,R,S,\theta)=\phi(W,\theta)+H\delta(W,\theta).$ (3.1)
Our work below shows that $\delta(W,\theta)$ parameterizes estimator variance due to assignment, while $\phi(W,\theta)$ parameterizes variance due to random sampling. In what follows, we work directly with this expansion.
Example 3.3 (SATE).
Continuing Example 2.6 above, let $\bar{Y}=(1-p)Y(1)+pY(0)$, a convex combination that summarizes each unit’s potential outcome level. Then for the score $g(D,Y,\theta)=HY-\theta$, we have $\delta(W,\theta)=\bar{Y}$ and
$E_{n}[H_{i}\delta(W_{i})]=\bar{\delta}_{1}-\bar{\delta}_{0}=\frac{\operatorname{Cov}_{n}(D_{i},\bar{Y}_{i})}{\operatorname{Var}_{n}(D_{i})}.$
Intuitively, $E_{n}[H_{i}\delta(W_{i})]$ from Equation 3.1 isolates the estimator variance due to chance correlations between the assignments $D_{i}$ and outcome levels $\bar{Y}_{i}$. By contrast, $\phi(W,\theta)=Y(1)-Y(0)-\theta$ does not depend on assignments $D_{i}$, and $\operatorname{Var}(\phi(W,\theta))=\operatorname{Var}(Y(1)-Y(0))$ isolates the estimator variance due to random sampling of treatment effects.
3.1 Finite Population Estimand
Our first theorem studies GMM estimation of the finite population estimand $\theta_{n}$ solving $E_{n}[\phi(W_{i},\theta_{n})]=0$. We study the superpopulation estimand $\theta_{0}$ in Corollary 3.8 below. To state the theorem, define the GMM linearization matrix $\Pi=-(G^{\prime}MG)^{-1}G^{\prime}M\in\mathbb{R}^{d_{\theta}\times d_{g}}$ and denote $v_{D}=\operatorname{Var}(D)=p-p^{2}$.
Theorem 3.4 (GMM).
Suppose $D_{1:n}$ as in Definition 2.1. Require Assumption 3.1, 3.2. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R_{A}$, independent RV’s with
$V_{\delta}=v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\Gamma_{0}^{\prime}h|\psi)]=\min_{\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\Gamma^{\prime}h|\psi)].$
The residual term $R_{A}\sim\Gamma_{0}^{\prime}Z_{h}\,|\,Z_{h}\in A$ for $Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)])$.
Note that variance matrix $V_{\delta}\in\mathbb{R}^{d_{\theta}\times d_{\theta}}$ and the minimum is interpreted in the positive semidefinite [^14] In particular, we say $V_{\delta}=\min_{\Gamma}V(\Gamma)$ if $V_{\delta}\preceq V(\Gamma)$ for all $\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}$. sense. Theorem 3.4 shows that $\sqrt{n}(\widehat{\theta}-\theta_{n})$ is asymptotically distributed as an independent sum of a normal $\mathcal{N}(0,V_{\delta})$ and truncated normal $R_{A}$. The normal term $\mathcal{N}(0,V_{\delta})$ only depends on the “treatment assignment” component of the influence function, $\Pi\delta(W,\theta_{0})$. The variance is attenuated nonparametrically by the stratification variables $\psi$ and linearly by rerandomization covariates $h$. The truncated normal term
$R_{A}\sim\Gamma_{0}^{\prime}Z_{h}\,|\,Z_{h}\in A,\quad\;Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)]).$
[…]
To isolate the precision gains due to rerandomization, the following corollary specializes Theorem 3.4 to the case of stratification without rerandomization, as well as complete randomization, as defined in Examples 2.3 and 2.4.
Corollary 3.5 (Pure Stratification).
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.1. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V)$ with $V=v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})|\psi)]$. In particular, if $D_{1:n}\sim\operatorname{CR}(p)$ then $V=v_{D}^{-1}\operatorname{Var}(\Pi\delta(W,\theta_{0}))$.
Remark 3.6 (Efficiency Gain).
Corollary 3.5 highlights how fine stratification reduces the variance of GMM estimation to $V=v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})|\psi)]\leq v_{D}^{-1}\operatorname{Var}(\delta(W,\theta_{0}))$, a nonparametric improvement. Rerandomization as in Definition 2.1 provides a further linear variance reduction to $V_{\delta}=\min_{\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\Gamma^{\prime}h|\psi)]$, up to the residual imbalance term $R_{A}$.
Remark 3.7 (Design-Based Asymptotics).
Our results above show that $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R_{A}$, conditional [^15] See Proposition 8.11 in the appendix for a formal statement. on $W_{1:n}=(R_{i},S_{i}(1),S_{i}(0))_{i=1}^{n}$. This result is “design-based” in the sense that the variance in the limiting distribution arises solely due to randomness of the treatment assignments $D_{1:n}$. However, we impose structure on the sequence of populations $W_{1:n}$ ex-ante, assuming each population is drawn from a fixed measure, $W_{i}\sim F$. This allows us to provide intuitive, closed form variance expressions and connect our results with the literature on GMM and partially linear adjustment. By contrast, the finite populations model often used in the statistics literature (e.g. Li et al., (2018)) begins with an arbitrary sequence of finite populations, imposing the minimal structure needed for certain moments to converge ex-post. It may be possible to extend our results to this setting, but we leave this to future work.
3.2 Superpopulation Estimand
The next result extends Theorem 3.4 to the superpopulation estimand $\theta_{0}$, which uniquely solves $E[\phi(W,\theta_{0})]=0$.
Corollary 3.8 (Superpopulation Estimand).
Suppose $D_{1:n}$ is as in Definition 2.1. Require Assumption 3.1, 3.2.
  1. (a)
    We have $\sqrt{n}(\widehat{\theta}-\theta_{0})\Rightarrow\mathcal{N}(0,V_{\phi})+\mathcal{N}(0,V_{\delta})+R_{A}$, independent RV’s with $V_{\phi}=\operatorname{Var}(\Pi\phi(W,\theta_{0}))$ and $V_{\delta}$, $R_{A}$ exactly as in Theorem 3.4.
  2. (b)
    (Pure Stratification). If $A=\mathbb{R}^{d_{h}}$, this is $\sqrt{n}(\widehat{\theta}-\theta_{0})\Rightarrow\mathcal{N}(0,V)$ with
    $V=\operatorname{Var}(\Pi\phi(W,\theta_{0}))+v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})|\psi)].$
Comparing Corollary 3.8 with the results above, we see that targeting $\theta_{0}$ instead of $\theta_{n}$ adds an extra independent Gaussian term $\mathcal{N}(0,V_{\phi})$ to the asymptotic distribution. Intuitively, $V_{\phi}$ arises due to iid random sampling of $\Pi\phi(W,\theta_{0})$. Notice that stratification and rerandomization only affect the assignment influence function component $\Pi\delta(W,\theta_{0})$, while the sampling influence component $\Pi\phi(W,\theta_{0})$ is irreducible. For pure stratification, Bai et al., (2024) were the first to derive an analogue of part (b) of Corollary 3.8 in the exactly identified case, under different GMM regularity conditions than we use here.
Example 3.9 (SATE).
Continuing Example 2.6, we had $\phi(W,\theta)=Y(1)-Y(0)-\theta$, so $G=1$ and $\Pi=1$. As above, $\delta(W,\theta)=(1-p)Y(1)+pY(0)\equiv\bar{Y}$. The GMM estimator $\widehat{\theta}=\bar{Y}_{1}-\bar{Y}_{0}$ is just difference of means. Then by Theorem 3.4 and Corollary 3.8, we have $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R_{A}$ and $\sqrt{n}(\widehat{\theta}-\operatorname{ATE})\Rightarrow\mathcal{N}(0,V_{\phi}+V_{\delta})+R_{A}$ with
$V_{\phi}=\operatorname{Var}(Y(1)-Y(0))\quad\quad V_{\delta}=\min_{\gamma\in\mathbb{R}^{d_{h}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}-\gamma^{\prime}h|\psi)].$ (3.2)
The term $V_{\phi}$, which only appears when estimating the superpopulation estimand $\theta_{0}$, reflects sampling variance due to treatment effect heterogeneity. The term $V_{\delta}$ is the variance due to random assignment, caused by random in-sample correlations between treatments $D$ and outcome levels $\bar{Y}$. Covariate-adaptive randomization and adjustment can be used to reduce $V_{\delta}$, while $V_{\phi}$ is an irreducible sampling variance.
Remark 3.10.
Wang et al., (2021) study SATE estimation under stratified rerandomization in the sequence of finite populations framework. Relative to Wang et al., (2021), by imposing the tight-matching condition 2.1 we are able to derive a simple closed form for the asymptotic variance in terms of the measure $W\sim F$, showing an equivalence with partially linear regression adjustment.
Example 3.11 (Treatment Effect Heterogeneity).
Continuing Example 2.7, consider the case with perfect compliance $D=Z$ and $f(X,\theta)=X^{\prime}\theta$. Then we can use the slightly modified score $g(D,X,Y,\theta)=(HY-X^{\prime}\theta)X$. For $\tau=Y(1)-Y(0)$, the parameters
$\theta_{n}=\operatorname*{argmin}_{\theta}E_{n}[(\tau_{i}-X_{i}^{\prime}\theta)^{2}],\quad\quad\theta_{0}=\operatorname*{argmin}_{\theta}E[(\tau-X^{\prime}\theta)^{2}].$
We have $\phi(W,\theta_{0})=(\tau-X^{\prime}\theta_{0})X$, where $e=\tau-X^{\prime}\theta_{0}$ are treatment effect prediction errors. For $\delta(W,\theta_{0})=\bar{Y}X$ and $\Pi=E[XX^{\prime}]^{-1}$, the variance matrices above are
$V_{\phi}=E[XX^{\prime}]^{-1}E[e^{2}XX^{\prime}]E[XX^{\prime}]^{-1}\quad\quad V_{\delta}=\min_{\Gamma\in\mathbb{R}^{d_{h}\times d_{x}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}\Pi X-\Gamma^{\prime}h|\psi)].$
The expression for $V_{\delta}$ shows that if we want to precisely estimate treatment effect heterogeneity, it is important to balance not only the variables that predict outcome levels $\bar{Y}$, but also their interactions with the heterogeneity variable.
3.3 Equivalence with Partially Linear Adjustment
Example 3.9 showed that, up to the rerandomization imbalance $R_{A}$, the unadjusted estimator $\widehat{\theta}=\bar{Y}_{1}-\bar{Y}_{0}$ has asymptotic variance given by the residuals of a partially linear regression of $\bar{Y}$ on $\psi$ and $h$:
$V_{\delta}=\min_{\gamma\in\mathbb{R}^{d_{h}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}-\gamma^{\prime}h|\psi)]=\min_{\begin{subarray}{c}\gamma\in\mathbb{R}^{d_{h}}\\ t\in L_{2}(\psi)\end{subarray}}v_{D}^{-1}\operatorname{Var}(\bar{Y}-\gamma^{\prime}h-t(\psi)).$ (3.3)
More generally, Theorem 3.4 shows that under stratified rerandomization designs, the usual GMM estimator $\widehat{\theta}$ behaves like semiparametrically adjusted GMM. Formally, let $\mathcal{L}(\psi)=L_{2}^{d_{\theta}}(\psi)$ be the $d_{\theta}$-fold Cartesian product of $L_{2}(\psi)$. Then $V_{\delta}$ in Theorem 3.4 is the variance of the residuals of the influence function $\Pi\delta(W,\theta_{0})$ in a partially linear regression on $\psi$ and $h$:
$V_{\delta}=\min_{\begin{subarray}{c}\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}\\ t\in\mathcal{L}(\psi)\end{subarray}}v_{D}^{-1}\operatorname{Var}\left(\Pi\delta(W,\theta_{0})-\Gamma^{\prime}h-t(\psi)\right).$ (3.4)
Intuitively, stratified rerandomization does partially linear regression adjustment “by design,” providing nonparametric control over $\psi$ and linear control over $h$. For a more explicit equivalence statement, define $m(\psi,h)=\Gamma_{0}^{\prime}h+t_{0}(\psi)$ to be the partially linear function achieving the optimum in Equation 3.4. Define the oracle semiparametrically adjusted GMM estimator
$\widehat{\theta}^{*}=\widehat{\theta}-E_{n}[H_{i}m(\psi_{i},h_{i})].$ (3.5)
For example, for the $\operatorname{SATE}$ estimation problem one can show that $\widehat{\theta}^{*}$ is just an oracle version of the usual augmented inverse propensity weighting (AIPW) estimator (Robins and Rotnitzky, (1995)), with partially linear regression models in each arm. [^16] Feasible partially linear adjustment in an iid mean estimation problem with missing data was studied in Wang et al., (2004). See also the related semiparametric adjustment for GMM parameters in Graham, (2011).
Theorem 3.12 (Partially Linear Adjustment).
Suppose that $D_{1:n}\sim\operatorname{CR}(p)$. The oracle partially linearly adjusted GMM estimator $\sqrt{n}(\widehat{\theta}^{*}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})$, with variance $V_{\delta}$ as defined in Theorem 3.4.
Without stratification or rerandomization, we require ex-post semiparametric adjustment to achieve $V_{\delta}$. Under stratified rerandomization, however, the simple GMM estimator $\widehat{\theta}$ automatically achieves $V_{\delta}$, up to the imbalance term $R_{A}$.
4 Nonlinear Rerandomization
In this section, we study various nonlinear rerandomization criteria, showing that in many cases they are asymptotically equivalent to linear rerandomization as in Definition 2.1, with an implicit choice of rerandomization covariates and acceptance region. This shows that our asymptotics and inference methods apply to a broad class of asymptotically linear rerandomization schemes.
4.1 GMM Rerandomization
First, we generalize the imbalance metric $\mathcal{I}_{n}$ introduced in Definition 2.1, allowing rejection of a treatment allocation $D_{1:n}$ based on potentially nonlinear features of the in-sample distribution of treatments and covariates $(D_{i},X_{i})_{i=1}^{n}$. To define the nonlinear imbalance metric, let $m(X_{i},\beta)$ be a score function and define the within-arm GMM estimators $\widehat{\beta}_{1}$ and $\widehat{\beta}_{0}$ by
$\widehat{\beta}_{1}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{d_{\beta}}}|E_{n}[D_{i}m(X_{i},\beta)]|_{2}^{2},\quad\;\widehat{\beta}_{0}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{d_{\beta}}}|E_{n}[(1-D_{i})m(X_{i},\beta)]|_{2}^{2}.$ (4.1)
Definition 4.1 (GMM Rerandomization).
Define $\mathcal{I}_{n}^{m}=\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})$ as above, where $m(X,\beta)$ is a score satisfying Assumption 3.2. Suppose $d_{\beta}=d_{m}$ (exact identification) and let $A$ a symmetric acceptance region. Do the following: (1) form strata as in Definition 2.1. (2) Draw $D_{1:n}$ by stratified randomization. (3) If imbalance $\mathcal{I}_{n}^{m}=\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\in A$, accept. Otherwise, repeat from (2).
Observe that if $m(X_{i},\beta)=X_{i}-\beta$, then $\widehat{\beta}_{d}=\bar{X}_{d}$ and $\mathcal{I}_{n}^{m}=\mathcal{I}_{n}$, so linear rerandomization is a special case. Intuitively, the generalization $\mathcal{I}_{n}^{m}$ allows us to randomize until possibly nonlinear features of the covariates are balanced between the treatment and control groups.
Example 4.2 (Density Rerandomization).
Let $f(X,\beta)$ be a parametric density model for the covariates $X$, which may be misspecified. Consider forming likelihood estimators $\widehat{\beta}_{1}\in\operatorname*{argmax}_{\beta}E_{n}[D_{i}\log f(X_{i},\beta)]$ and $\widehat{\beta}_{0}\in\operatorname*{argmax}_{\beta}E_{n}[(1-D_{i})\log f(X_{i},\beta)]$, then rerandomizing until the imbalance measure $\sqrt{n}|\widehat{\beta}_{1}-\widehat{\beta}_{0}|_{2}\leq\epsilon$. Under suitable regularity conditions, $\widehat{\beta}_{d}$ are GMM estimators as in Equation 4.1 with score function $m(X_{i},\beta)=\nabla_{\beta}f(X_{i},\beta)$, so this procedure is a GMM rerandomization with acceptance region $A=\{x:|x|_{2}\leq\epsilon\}$.
Let $\beta^{*}$ be the unique solution to $E[m(X,\beta^{*})]=0$ and define $G_{m}=E[(\partial/\partial\beta^{\prime})m(X_{i},\beta^{*})]$. Our next result shows that GMM rerandomization with acceptance criterion $\mathcal{I}_{n}^{m}\in A$ is equivalent to linear rerandomization (Definition 2.1) with an implicit choice of rerandomization covariates $h_{i}=m(X_{i},\beta^{*})$ and linearly transformed acceptance region.
Theorem 4.3 (GMM Rerandomization).
Suppose $D_{1:n}$ is as in Definition 4.1 and Assumption 3.2 holds. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R$, independent RV’s with
$V_{\delta}=\min_{\eta\in\mathbb{R}^{d_{\beta}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\eta^{\prime}m(X_{i},\beta^{*})|\psi)].$
The residual $R\sim\eta_{0}^{\prime}Z_{m}\,|\,Z_{m}\in G_{m}A$ for $Z_{m}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(m(X_{i},\beta^{*})|\psi)])$.
Theorem 4.3 shows that by rerandomizing until $\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\in A$, we implicitly balance the influence function for the difference of GMM estimators in Equation 4.1, which depends both on $m(X_{i},\beta^{*})$ and the Jacobian matrix $G_{m}$.
This suggests an equivalent, but computationally simpler design with only one round of nonlinear estimation. In particular, let $\widehat{\beta}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{d_{\beta}}}|E_{n}[m(X_{i},\beta)]|_{2}^{2}$ and set rerandomization covariates $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$, rerandomizing until $\sqrt{n}E_{n}[H_{i}\widehat{h}_{i}]\in G_{m}A$. This design is asymptotically equivalent to Definition 4.1, as shown in the next result.
Corollary 4.4.
Suppose Assumption 3.1, 3.2 hold and let score $m(X,\beta)$ as in Definition 4.1. Let $D_{1:n}$ be rerandomized as in Definition 2.1 with $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$ and acceptance region $G_{m}A$. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R$, with both variables identical to those in Theorem 4.3.
Corollary 4.4 shows that we can achieve the same effect as Definition 4.1 with a computationally simpler linear rerandomization that balances the estimated covariate $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$. Next, we introduce a propensity score based approach that also attempts to balance nonlinear covariate features.
4.2 Propensity Rerandomization
To motivate a propensity based rerandomization procedure, note that under stratified randomization we have $E[D_{i}|W]=p$ for all units. In finite samples, however, the realized propensity $\widehat{p}(S)=E_{n}[D_{i}|X_{i}\in S]$ may be significantly different from $p$ in certain regions $S\subseteq\mathbb{R}^{d_{X}}$ of the covariate space. This implies that covariates are predictive of treatment assignments post-randomization, a form of “in-sample confounding.” To enforce balance, we could, for instance, reject allocations where $|\widehat{p}(S)-p|>\epsilon$ for some collection of sets $S$. To make this idea tractable without fully discretizing, consider a parametric propensity model $p(X,\beta)=L(X^{\prime}\beta)$ and define the MLE estimator
$\widehat{\beta}\in\operatorname*{argmax}_{\beta\in\mathbb{R}^{d_{\beta}}}E_{n}[D_{i}\log L(X_{i}^{\prime}\beta)+(1-D_{i})\log(1-L(X_{i}^{\prime}\beta))].$ (4.2)
We can measure the average gap between the estimated and true propensity score using the imbalance metric
$\mathcal{J}_{n}=nE_{n}[(p-L(X_{i}^{\prime}\widehat{\beta}))^{2}].$ (4.3)
Intuitively, if $\mathcal{J}_{n}$ is large, then the covariates $X$ have power to predict treatment status in some parts of the covariate space. To avoid this, we propose rerandomizing until the imbalance metric $\mathcal{J}_{n}$ is below a threshold:
Definition 4.5 (Propensity Rerandomization).
Do the following: (1) form strata as in Definition 2.1. (2) Draw $D_{1:n}$ and estimate the propensity model in Equation 4.2. (3) If imbalance $\mathcal{J}_{n}\leq\epsilon$, accept. Otherwise, repeat from (2).
Before we state the main result, we require some regularity conditions
Assumption 4.6.
Impose the following conditions.
  1. (a)
    Let $L$ be twice differentiable, with $|L^{\prime}|_{\infty},|L^{\prime\prime}|_{\infty}<\infty$. For each $p\in(0,1)$, there is a unique $c$ with $L(c)=p$. Also, $|L^{\prime}(c)|>0$.
  2. (b)
    The score $m(D_{i},X_{i},\beta)=D_{i}\frac{L^{\prime}(X_{i}^{\prime}\beta)X_{i}}{L(X_{i}^{\prime}\beta)}-(1-D_{i})\frac{L^{\prime}(X_{i}^{\prime}\beta)X_{i}}{1-L(X_{i}^{\prime}\beta)}$ satisfies condition 3.2. The solution to Equation 4.2 exists and is unique.
  3. (c)
    Covariates $X=(1,h)$ for $E[|h|^{2}_{2}]<\infty$. Also, $E[\operatorname{Var}(h|\psi)]$, $E[XX^{\prime}]$ are full rank.
Condition (a) is satisfied by the logit and probit link functions. Our next result shows that propensity rerandomization as in Definition 4.5 is equivalent to a simpler linear rerandomization design, with an implicit choice of ellipsoidal acceptance region.
Theorem 4.7 (Propensity Rerandomization).
Suppose $D_{1:n}$ is as in Definition 4.5. Require Assumptions 3.2, 4.6. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R$.
$V_{\delta}=\min_{\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\Gamma^{\prime}h|\psi)].$
The residual $R\sim\Gamma_{0}^{\prime}Z_{h}\,|\,Z_{h}^{\prime}\operatorname{Var}(h)^{-1}Z_{h}\leq\epsilon$ for $Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)])$ and $\Gamma_{0}$ optimal in the equation above.
Theorem 4.7 shows that for any sufficiently regular link function, propensity rerandomization is asymptotically equivalent to the simpler quadratic rerandomization design in Example 2.2, with acceptance criterion $n(\bar{h}_{1}-\bar{h}_{0})^{\prime}\operatorname{Var}_{n}(h_{i})^{-1}(\bar{h}_{1}-\bar{h}_{0})\leq\epsilon$. Equivalently, propensity rerandomization behaves like linear rerandomization with $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ and ellipsoidal acceptance region $A=\operatorname{Var}(h)^{1/2}B(0,\epsilon)$. [^17] A related result was found by Ding and Zhao, (2024), who study rerandomizing until the p-value of a logistic regression coefficient is above a threshold.
Implicit Acceptance Regions. Both nonlinear designs in this section turned out to be equivalent to the standard rerandomization scheme in Definition 2.1, with a specific, implicit choice of rerandomization moments and acceptance region. However, note that the moments and acceptance region chosen by these procedures are entirely determined by the marginal covariate distribution. This implicit choice is not likely to be optimal, since the residual term $R_{A}\sim\Gamma_{0}^{\prime}Z_{h}|Z_{h}\in A$ depends not only on the covariates $Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)])$, but also on the partially linear coefficient $\Gamma_{0}$. This coefficient is determined by the joint distribution of the assignment influence function $\Pi\delta(W,\theta_{0})$ and covariates $(\psi,h)$. In the next section, we show how to use prior information about this joint distribution to optimize the acceptance region and bound the variance of $R_{A}$.
5 Optimizing Acceptance Regions
In this section, we study efficient choice of the rerandomization acceptance region $A\subseteq\mathbb{R}^{d_{h}}$. For simplicity and intuition, we first restrict to the case of estimating $\theta_{n}=\operatorname{SATE}$, generalizing in what follows.
Imbalance Decomposition. The difference of means estimator $\widehat{\theta}=E_{n}[Y_{i}(1)-Y_{i}(0)]+E_{n}[H_{i}\bar{Y}_{i}]=\theta_{n}+E_{n}[H_{i}\bar{Y}_{i}]$ for $H_{i}=(D_{i}-p)/(p-p^{2})$. Intuitively, the scaled errors $\sqrt{n}(\widehat{\theta}-\theta_{n})=\sqrt{n}E_{n}[H_{i}\bar{Y}_{i}]$ are driven by imbalances in the outcome levels $\bar{Y}$ between treatment arms. The previous section showed that if covariates $h$ are predictive of $\bar{Y}$, we can reduce these imbalances by rerandomizing until $\sqrt{n}E_{n}[H_{i}h_{i}]=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})\in A$. To study the role of the acceptance region $A$, let $(\gamma_{0},t_{0})$ be solutions to the partially linear prediction problem in Equation 3.3 and consider the expansion [^18] This is without loss of generality. Note that we do not impose well-specification $E[e|\psi,h]=0$.
$\bar{Y}=\gamma_{0}^{\prime}h+t_{0}(\psi)+e,\quad\;E[e|\psi]=0,\;E[eh]=0.$
We use this to decompose the imbalance in outcome levels $\bar{Y}$ into imbalances in the covariates $\psi$ and $h$ and residuals $e$. In particular, we can now write the imbalance decomposition
$E_{n}[H_{i}\bar{Y}_{i}]=E_{n}[H_{i}t_{0}(\psi_{i})]+\gamma_{0}^{\prime}E_{n}[H_{i}h_{i}]+E_{n}[H_{i}e_{i}]\equiv I_{1}+I_{2}+I_{3}.$
The analysis in Section 3 showed that for any acceptance region $A\subseteq\mathbb{R}^{d_{h}}$:
  1. (1)
[…]
  2. (2)
    The components $I_{2}+I_{3}=\gamma_{0}^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]+\sqrt{n}E_{n}[H_{i}e_{i}]\Rightarrow R_{A}+v_{D}^{-1/2}\mathcal{N}(0,\operatorname{Var}(e))$ are asymptotically independent, with $\operatorname{Var}(e)$ not depending on $A$.
In particular, it suffices to choose $A$ to minimize the component $I_{2}=\gamma_{0}^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]$. This suggests an oracle acceptance criterion, rerandomizing until $|\gamma_{0}^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]|\leq\epsilon$, with acceptance region $A=\{a:|\gamma_{0}^{\prime}a|\leq\epsilon\}$. However, this acceptance region is infeasible since $\gamma_{0}$ is unknown at design-time. Instead, we take a minimax approach, allowing the researcher to incorporate prior information about $\gamma_{0}$.
5.1 Minimax Rerandomization
Suppose that we know $\gamma_{0}\in B$ for some prior information set $B\subseteq\mathbb{R}^{d_{h}}$. Fix $\epsilon>0$ and consider a “minimax” style acceptance criterion, rerandomizing the treatments $D_{1:n}$ until
$\sup_{\gamma\in B}|\gamma^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]|\leq\epsilon.$ (5.1)
Note that the function $f_{B}(x)=\sup_{\gamma\in B}|\gamma^{\prime}x|$ is convex, so we can also interpret this as a convex imbalance penalty, rerandomizing until $f_{B}(\mathcal{I}_{n})\leq\epsilon$ for imbalance metric $\mathcal{I}_{n}=\sqrt{n}E_{n}[H_{i}h_{i}]$, generalizing the quadratic penalty in Example 2.2. Our first result shows that this minimax design is of the form studied in the Section 3, characterizing the acceptance region induced by this convex penalty.
Theorem 5.1.
The following hold:
  1. (a)
    (Rerandomization). The acceptance criterion $\sup_{\gamma\in B}|\gamma^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]|\leq\epsilon$ $\iff$ $\sqrt{n}E_{n}[H_{i}h_{i}]\in A$ for $A=\epsilon B^{\circ}$ with $B^{\circ}=\{a:\sup_{\gamma\in B}|\gamma^{\prime}a|\leq 1\}\subseteq\mathbb{R}^{d_{h}}$. [^19] Note that for a set $S\subseteq\mathbb{R}^{d}$, we have $\epsilon S=\{\epsilon s:s\in S\}$.
  2. (b)
    (Acceptance Region). $A=\epsilon B^{\circ}$ is symmetric and convex. If $B$ is bounded, then $A$ is closed and has non-empty interior. If $B$ is open, then $A$ is bounded.
  3. (c)
    (Well-specification). If $\gamma_{0}\in B$, then $\operatorname{Var}(R_{A})\leq\epsilon^{2}$.
Part (a) of Theorem 5.1 shows that the rerandomization criterion is of the form studied in Definition 2.1, with acceptance region $A=\epsilon B^{\circ}$. Part (b) shows that $A$ is always symmetric and convex. In particular, the asymptotic distribution of $\widehat{\theta}$ is centered at zero. The set $B^{\circ}$ is known as the absolute polar of $B$, e.g. see Aliprantis and Border, (2006). Part (c) of the theorem shows that if the prior information set $B$ contains the true coefficient $\gamma_{0}$, then $\operatorname{Var}(R_{A})\leq\epsilon^{2}$. Then by independence, the asymptotic variance is within $\epsilon^{2}$ of the optimal partially linear variance. If $\gamma_{0}\not\in B$ (misspecification), then possibly $\operatorname{Var}(R_{A})>\epsilon^{2}$. However, note that misspecification does not affect our inference methods, which allow for general acceptance regions $A$.
Remark 5.2 (Acceptance Probability).
Note that the asymptotic acceptance probability $a(\epsilon)=P(Z_{h}\in\epsilon B^{\circ})$ has $a(\epsilon)\to 0$ as $\epsilon\to 0$ and is monotonically increasing. For $B$ bounded, the theorem shows that $B^{\circ}$ has non-empty interior. In this case, as $\epsilon\to\infty$ we have $\epsilon B^{\circ}\uparrow\mathbb{R}^{d_{h}}$ so $a(\epsilon)\to 1$. This shows that, at least in large samples, we can choose $\epsilon$ to achieve any desired acceptance probability $P(Z_{h}\in A)\in(0,1)$. Under well-specification, any such choice of $\epsilon$ comes with a variance guarantee provided by the theorem.
5.2 Specifying Prior Information
Without pilot data, we are left to introspection to choose the prior information set $B$. Recall that $\gamma_{0}$ is the coefficient from the partially linear projection $\bar{Y}=\gamma_{0}^{\prime}h+t_{0}(\psi)+e$. Intuitively, $\gamma_{0}$ parameterizes how much we expect the average outcome level to change given a unit change in $h$, holding $\psi$ fixed. If the partially linear model happens to be well-specified, then $\gamma_{0}=\nabla_{b}E[\bar{Y}|\psi,h=b]$. If $t_{0}(\psi)=t^{\prime}\psi$ happens to be linear, then $\bar{Y}=c+\gamma_{0}^{\prime}h+t^{\prime}\psi+e$ and $\gamma_{0}$ is just an OLS coefficient. The following examples provide some reasonable prior information specifications and their associated acceptance regions. These examples rely on a general characterization of acceptance regions in Lemma 5.5 below.
Example 5.3 (Rectangle).
One natural way to specify prior information is to assume $\gamma_{0j}\in[l_{j},u_{j}]$ for each $1\leq j\leq d_{h}$, equivalent to setting $B=\prod_{j=1}^{d_{h}}[l_{j},u_{j}]$. This allows sign constraints, e.g. $0\leq\gamma_{0j}\leq m$ for some $j$ and $-m\leq\gamma_{0j}\leq 0$ for others. Lemma 5.5 below shows that if $B=\prod_{j=1}^{d_{h}}[l_{j},u_{j}]$, then $A=\epsilon B^{\circ}=\{a:|a^{\prime}l+a^{\prime}u|+\sum_{j}|a_{j}|u_{j}-|a_{j}|l_{j}\leq 2\epsilon\}$, where $l=(l_{j})_{j}$ and $u=(u_{j})_{j}$. An example is shown in Figure 1. Note that the acceptance region $A$ is conservative in directions aligned with the prior information set $B=[1,2]\times[1,3/2]$, guarding against covariate imbalances that are aligned with adverse coefficient values $\gamma_{0}\in B$. $A$ is more lenient in directions approximately orthogonal to $B$.
Example 5.4 (Ellipse).
Another natural specification is to guess $\gamma_{0}\approx\bar{\gamma}$, setting $B=\bar{\gamma}+B_{2}(0,m)$, for an uncertainty parameter $m$. By the characterization in Lemma 5.5 below, $A=\epsilon B^{\circ}=\{a:|a^{\prime}\bar{\gamma}|+m|a|_{2}\leq\epsilon\}$. More generally, if $B=\bar{\gamma}+\Sigma B_{2}(0,1)$ for a positive-definite matrix $\Sigma$, the lemma shows that $A=\epsilon B^{\circ}=\{a:|a^{\prime}\bar{\gamma}|+|\Sigma a|_{2}\leq\epsilon\}$. One natural application of this specification is when $B$ is a Wald confidence region constructed using pilot data, as discussed below. An example is shown in Figure 1.
Refer to caption Refer to caption
Figure 1: Prior information $B$ and $A=\epsilon B^{\circ}$ for Examples 5.3 and 5.4.
More generally, the following lemma provides a useful characterization of the acceptance region $A=\epsilon B^{\circ}$ from Theorem 5.1 for a large family of prior information set specifications. To state the lemma, recall that $|x|_{p}=(\sum_{j}|x_{j}|^{p})^{1/p}$ for $p\in[1,\infty)$ and $|x|_{\infty}=\max_{j}|x_{j}|$. For $p\in[1,\infty]$, denote $B_{p}(0,1)=\{a:|a|_{p}\leq 1\}$.
Lemma 5.5 (Acceptance Regions).
For $p\in[1,\infty]$, let $11/p+1/q=$ 1, setting $q=1$ if $p=\infty$ and vice-versa. Suppose $B=x+\Sigma B_{p}(0,1)$, for $x\in\mathbb{R}^{d_{h}}$ and $\Sigma$ invertible. Then $A=\epsilon B^{\circ}=\{a:|a^{\prime}x|+|\Sigma^{\prime}a|_{q}\leq\epsilon\}$.
5.3 Using Pilot Data
Next, we discuss an alternative strategy that uses pilot data to specify the set $B$. Suppose we have access to $\mathcal{D}_{pilot}\perp\!\!\!\perp(W_{1:n},D_{1:n})$ of size $m$. Suppose $\sqrt{m}(\widehat{\gamma}_{pilot}-\gamma_{0})\approx\mathcal{N}(0,\widehat{\Sigma}_{pilot})$ for some estimator $\widehat{\gamma}_{pilot}$, discussed below. Consider forming the Wald region $\widehat{B}_{pilot}=\{\gamma:m(\widehat{\gamma}_{pilot}-\gamma)^{\prime}\widehat{\Sigma}_{pilot}^{-1}(\widehat{\gamma}_{pilot}-\gamma)\leq c_{\alpha}\}$ using critical value $P(\chi^{2}_{d_{h}}\leq c_{\alpha})=1-\alpha$ for $\alpha\in(0,1)$. Equivalently, one can write the Wald region as
$\widehat{B}_{pilot}=\widehat{\gamma}+c_{\alpha}^{1/2}m^{-1/2}\cdot\widehat{\Sigma}_{pilot}^{1/2}B_{2}(0,1).$ (5.2)
Using $\widehat{B}_{pilot}$ as a prior information set, by Example 5.4 we have acceptance region
$\widehat{A}_{pilot}=\epsilon\widehat{B}_{pilot}^{\circ}=\{a:|a^{\prime}\widehat{\gamma}_{pilot}|+m^{-1/2}c_{\alpha}^{1/2}|\widehat{\Sigma}^{1/2}a|_{2}\leq\epsilon\}.$ (5.3)
Note that the acceptance region $\widehat{A}_{pilot}$ grows with the pilot size $m$. This reflects smaller uncertainty about the true parameter $\gamma_{0}$, and thus less adversarial worst case imbalance $\sup_{\gamma\in\widehat{B}_{pilot}}|\gamma^{\prime}E_{n}[H_{i}h_{i}]|$. Conversely, $\widehat{A}_{pilot}$ shrinks as the confidence parameter $\alpha$ and the scale of the variance estimate $\widehat{\Sigma}_{pilot}$ increases, reflecting greater uncertainty and a more conservative approach to covariate balances. Our next result shows that rerandomization with acceptance region $\widehat{A}_{pilot}$ controls the variance of the imbalance $R_{A}=\gamma_{0}^{\prime}Z|Z\in\widehat{A}_{pilot}$ with high probability marginally over the realizations of the pilot data. The result is an immediate consequence of Theorem 3.4 and Theorem 5.1.
Corollary 5.6 (Pilot Data).
Suppose $P(\gamma_{0}\in\widehat{B}_{pilot})\geq 1-\alpha$, for $\mathcal{D}_{pilot}\perp\!\!\!\perp(W_{1:n},D_{1:n})$. Let $D_{1:n}$ as in Definition 2.1 with $A=\widehat{A}_{pilot}=\epsilon\widehat{B}_{pilot}^{\circ}$, then if Assumptions 3.1, 3.2 hold, then $\sqrt{n}(\widehat{\theta}-\theta_{n})|\mathcal{D}_{pilot}\Rightarrow v_{D}^{-1}\mathcal{N}(0,\operatorname{Var}(e))+R_{A}$ and $\operatorname{Var}(R_{A}|\mathcal{D}_{pilot})\leq\epsilon^{2}$ with probability $\geq 1-\alpha$.
Formally, the pilot estimate of $\gamma_{0}$ and Wald region could be constructed as in Robinson, (1988). In practice, a simple approach suggested by the theory is to let $\widehat{\gamma}_{pilot},\widehat{\Sigma}_{pilot}$ be point and variance estimators from the regression $Y_{T}\sim 1+h+\psi$, for the “tyranny of the minority” (Lin, (2013)) outcomes $Y_{T}=(1-p)DY/p+p(1-D)Y/(1-p)$, noting that $E[Y_{T}|W]=(1-p)Y(1)+pY(0)=\bar{Y}$.
General Parameters. For completeness, we extend the preceding work to general parameters $\theta_{n}$ as in Definition 2.5. Let $\Pi\delta(W,\theta_{0})$ be the assignment influence function. As in Equation 3.4, consider the partially linear decomposition
$\Pi\delta(W,\theta_{0})=\Gamma_{0}^{\prime}h+t_{0}(\psi)+e,\quad\;E[e|\psi]=0,\;E[eh]=0.$
Note that $e\in\mathbb{R}^{d_{\theta}}$ and $E[e|\psi]=0$ is interpreted componentwise. Consider prior information sets $B_{j}$ for each $\Gamma_{0}^{j}$ with $1\leq j\leq d_{\theta}$, where $\Gamma_{0}^{j}\in\mathbb{R}^{d_{h}}$ is the $j$th column of $\Gamma_{0}$. The final result of this section bounds the asymptotic imbalance term $R_{A}$ if all these prior information sets are well specified.
Theorem 5.7.
Let $D_{1:n}$ as in Definition 2.1 with $A=\cap_{j=1}^{d_{\theta}}\epsilon B^{\circ}_{j}$. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R_{A}$, as defined in Theorem 3.4. If $\Gamma_{0}^{j}\in B_{j}$ $\forall j$, then $\max_{j=1}^{d_{\theta}}\operatorname{Var}((R_{A})_{jj})\leq\epsilon^{2}$.
Note that by construction the conservative acceptance region $A=\cap_{j=1}^{d_{\theta}}\epsilon B^{\circ}_{j}$ is symmetric and convex.
6 Linear Adjustment
In this section, we study optimal linearly adjusted GMM estimation under stratified rerandomization. We show that this can be used to completely remove the impact of the acceptance region and imbalance term $R_{A}$ to first order, restoring asymptotic normality. This allows for standard t-statistic and Wald-test based inference on the parameters $\theta_{n}$ and $\theta_{0}$, provided in Section 7 below. Let $w$ denote the covariates used for ex-post adjustment and suppose $E[|w|_{2}^{2}]<\infty$.
Definition 6.1 (Adjusted GMM).
[…]
First, we extend Corollary 3.5 to provide asymptotics for the adjusted GMM estimator under pure stratification ($A=\mathbb{R}^{d_{h}}$).
Proposition 6.2 (Linear Adjustment).
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.2. Then we have $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V(\alpha))$ with $V(\alpha)=v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\alpha^{\prime}w|\psi)]$ and $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{0})\Rightarrow\mathcal{N}(0,V_{\phi}+V(\alpha))$.
A version of this result was given in Cytrynbaum, (2023) for the special case $\theta_{0}=\operatorname{ATE}$. Motivated by Proposition 6.2, we define the optimal linear adjustment coefficient as the minimizer of the asymptotic variance $V(\alpha)$, in the positive semidefinite sense.
Optimal Adjustment Coefficient. Define the coefficient
$\alpha_{0}\in\operatorname*{argmin}_{\alpha\in\mathbb{R}^{d_{w}\times d_{g}}}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\alpha^{\prime}w|\psi)].$ (6.1)
Note that if $w=h$ then $\alpha_{0}=\Gamma_{0}$, as in Theorem 3.4. If $E[\operatorname{Var}(w|\psi)]\succ 0$, then the unique minimizer of Equation 6.1 is the partially linear regression coefficient $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\Pi\delta(W,\theta_{0})|\psi)]$. The main result of this section shows that adjustment by a consistent estimate of $\alpha_{0}$ restores asymptotic normality.
Theorem 6.3 (Adjusted GMM).
[…]
$V_{\phi}=\operatorname{Var}(\Pi\phi(W,\theta_{0}))\quad\quad V_{\delta}^{a}=\min_{\alpha\in\mathbb{R}^{d_{w}\times d_{g}}}v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\alpha^{\prime}w|\psi)].$
Two-step Adjustment. For nonlinear models, the coefficient $\alpha_{0}$ may depend on the unknown parameter $\theta_{0}$. This suggests a two-step adjustment strategy, where we
  1. (1)
[…]
  2. (2)
    Report the adjusted estimator $\widehat{\theta}_{adj}=\widehat{\theta}-E_{n}[H_{i}\widehat{\alpha}^{\prime}w_{i}]$.
[…]
Theorem 6.4 (Feasible Adjustment).
Suppose $D_{1:n}$ is as in Definition 2.1. Require Assumption 3.1, 3.2. Assume that $E[\operatorname{Var}(w|\psi)]\succ 0$. Define $\widehat{\alpha}=v_{D}E_{n}[\check{w}_{i}\check{w}_{i}^{\prime}]^{-1}E_{n}[H_{i}\check{w}_{i}\widehat{g}_{i}^{\prime}]\widehat{\Pi}^{\prime}$. Then $\widehat{\alpha}=\alpha_{0}+o_{p}(1)$.
Remark 6.5 (One Step Adjustment).
In many cases, the optimal coefficient $\alpha_{0}$ may not depend on $\theta_{0}$ at all, allowing optimal adjustment to be done in one step. For instance, whenever $\delta(W,\theta)=u(\psi,\theta)+v(W)$ for some functions $u,v$, then $\alpha_{0}$ does not depend on $\theta_{0}$. This happens in Example 3.9, where we have $\delta(W,\theta)=(1-p)Y(1)+pY(0)=\bar{Y}$. The optimal adjustment coefficient is given by $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\bar{Y}|\psi)]$.
Theorem 6.3 that ex-post linear adjustment can be used to remove the non-Gaussian component of the asymptotic distribution that arises due to rerandomization. In the next section, we exploit this result to provide standard t-statistic based confidence intervals for $\theta_{n}$ and $\theta_{0}$.
7 Inference
In this section, we provide novel methods for inference on general causal parameters under stratified rerandomization designs. We make crucial use of asymptotic normality of the optimally adjusted estimator $\widehat{\theta}_{adj}$, shown in Theorem 6.3. For the superpopulation parameter $\theta_{0}$, we provide asymptotically exact inference methods. The asymptotic variance for estimating the finite population parameter $\theta_{n}$ is generally not identified. In this case, we provide conservative variance estimation that still reflects the precision gains due to stratification and rerandomization.
7.1 Asymptotically Exact Inference
To define our variance estimator, we begin with some definitions. Let $\mathcal{G}_{n}$ denote the set of groups constructed in Definition 2.1. For each $g\in\mathcal{G}_{n}$ define the centroid $\bar{\psi}_{g}=|g|^{-1}\sum_{i\in g}\psi_{i}$. Let $\nu:\mathcal{G}_{n}\to\mathcal{G}_{n}$ be a bijective matching between groups satisfying $\nu(g)\not=g$, $\nu^{2}=\operatorname{Id}$, and the homogeneity condition
$\frac{1}{n}\sum_{g\in\mathcal{G}_{n}}|\bar{\psi}_{g}-\bar{\psi}_{\nu(g)}|_{2}^{2}=o_{p}(1).$ (7.1)
[…]
$\widehat{v}_{1}$ =$n^{-1}\sum_{g\in\mathcal{G}^{\nu}_{n}}\frac{1}{a(g)-1}\sum_{i\not=j\in g}\widehat{m}_{i}\widehat{m}_{j}^{\prime}D_{i}D_{j}/p$
$\widehat{v}_{0}$ =$n^{-1}\sum_{g\in\mathcal{G}^{\nu}_{n}}\frac{1}{(k-a)(g)-1}\sum_{i\not=j\in g}\widehat{m}_{i}\widehat{m}_{j}^{\prime}(1-D_{i})(1-D_{j})/(1-p)$
$\widehat{v}_{10}$ $=n^{-1}\sum_{g\in\mathcal{G}_{n}}\frac{k}{a(k-a)}(g)\sum_{i,j\in g}\widehat{m}_{i}\widehat{m}_{j}^{\prime}D_{i}(1-D_{j}).$
Using these terms, construct the variance estimator
$\widehat{V}=\operatorname{Var}_{n}(\widehat{m}_{i})-v_{D}(\widehat{v}_{1}+\widehat{v}_{0}-\widehat{v}_{10}-\widehat{v}_{10}^{\prime}).$ (7.2)
We require a slight strengthening of our GMM assumptions 3.2.
Assumption 7.1.
There exists $\theta_{0}\in U\subseteq\Theta$ open s.t. $E[\sup_{\theta\in U}|\partial/\partial\theta^{\prime}g_{d}(W,\theta)|_{F}^{2}]<\infty$.
Under this condition, we can state our first inference result, showing consistent estimation of the asymptotic variance matrix in Theorem 6.3.
Theorem 7.2 (Inference).
[…]
By Theorem 6.3, $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{0})\Rightarrow N(0,V_{\phi}+V_{\delta}^{a})$. Then the variance estimation result above allows for joint inference on $\theta_{0}$ using e.g. standard Wald-test or t-statistic based confidence regions.
7.2 Inference on the Finite Population Parameter
In this section, we provide asymptotically conservative inference on linear contrasts of the finite population parameter $c^{\prime}\theta_{n}$.
As noted above, the asymptotic variance $V_{\delta}^{a}$ in Theorem 6.3 for estimating the finite population parameter $\theta_{n}$ is generically not identified. This happens because it depends on terms of the form $\operatorname{Var}(\delta|\psi)\propto\operatorname{Var}(g_{1}|\psi)+\operatorname{Var}(g_{0}|\psi)-2\operatorname{Cov}(g_{1},g_{0}|\psi)$, with $g_{d}=g(d,X,S(d),\theta_{0})$. However, $S(1)$ and $S(0)$ are never simultaneously observed (Neyman, (1990)), so $\operatorname{Cov}(g_{1},g_{0}|\psi)$ is generically not identified. We work with linear contrasts $c^{\prime}\theta_{n}$ since this allows us to tighten our upper bounds on the (non-identified) variance. To do so, let $\widehat{u}_{1}=E_{n}[\frac{D_{i}}{p}\widehat{m}_{i}\widehat{m}_{i}^{\prime}]-\widehat{v}_{1}$ and $\widehat{u}_{0}=E_{n}[\frac{1-D_{i}}{1-p}\widehat{m}_{i}\widehat{m}_{i}^{\prime}]-\widehat{v}_{0}$ using the estimator components above and consider the variance estimator
$\widehat{V}_{\delta}(c)=v_{D}([c^{\prime}\widehat{u}_{1}c]^{1/2}+[c^{\prime}\widehat{u}_{0}c]^{1/2})^{2}.$ (7.3)
By Theorem 6.3, we have $\sqrt{n}(c^{\prime}\widehat{\theta}_{adj}-c^{\prime}\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,c^{\prime}V_{\delta}^{a}c)$. Our next result shows how to consistently estimate an upper bound on this asymptotic variance.
Theorem 7.3 (Inference).
[…]
The variance upper bound $\bar{V}_{\delta}(c)\geq c^{\prime}(V_{\phi}+V_{\delta}^{a})c$, so the confidence intervals derived from this approach are always weakly shorter than those using the variance estimator in Equation 7.2. See Section 8.6 in the appendix for an explicit comparison. The upper bound $\bar{V}_{\delta}(c)$ incorporates the efficiency gains from stratification, rerandomization, and adjustment. However, this upper bound is generally not sharp (Aronow et al., (2014)). We leave sharp upper bounds on the asymptotic variance matrix $V_{\delta}^{a}$ to future work.